# HPVM-HDC: A Heterogeneous Programming System for Accelerating Hyperdimensional Computing

### v2

@@ 3.1. High-Level Stage Primitives
In addition to describing individual HDC operations, HDC++ enables describing three commonly used HDC algorithmic stages: encoding_loop, training_loop, inference_loop. Listing 2 illustrates expressing an HDC inference loop using the corresponding primitive. There are two advantages of providing such high-level primitives. First, it enables HPVM-HDC to target HDC accelerators, which provide coarse-grain algorithm acceleration, in contrast to acceleration of finer-grain HDC operations. Second, the use of loop intrinsics for the key algorithmic stages implemented by the HDC accelerators allows HPVM-HDC to elide data movements to and from the accelerators. Section  4.3 describes HPVM-HDC in detail. Crucially, these high-level stage primitives are composable with more granular primitives in HDC++—this allows us to express certain applications that map partially, but not fully to HDC accelerators, such as HD-Clustering. By writing HD-Clustering in HDC++, we are able to map the computationally intensive part of clustering, HDC inference, to HDC accelerators, while performing more ancillary tasks, such as cluster updates and the initial generation of the random projection matrix, on the CPU or GPU.
The algorithmic stage primitives in HDC++ each take as input an ”implementation” function. This function implements encoding, training, or inference using more granular HDC primitives, and is executed when targeting CPUs or GPUs, rather than HDC accelerators. This is because while HDC accelerators implement specific encoding, training, and inference algorithms, CPUs and GPUs can be programmed to implement a variety of algorithms—it is up to the application developer to choose concrete versions of these algorithms for these targets. The benefits of this approach are elaborated on in Section 4.3.
Listing 2: HDC++ example of an inference loop. QueriesHM are input vectors to classify, ClassHM contains representative hypervectors per class (see Section 2.1), InferenceResults is the destination pointer to store the results in, and InferenceFunction is a function implementing the inference algorithm using HDC primitives.
@@ 4.1. The HPVM-HDC IR
HDC++ application code is compiled into an IR that contains HDC primitives as well as generic parallel task and loop primitives. The HDC primitives are LLVM intrinsics which represent the same operations as the HDC++ primitives listed in Table 1. The generic parallel task and loop primitives come from HPVM IR, which represents code using LLVM IR with additional annotations and an auxiliary hierarchical dataflow graph representing both data and task parallelism. The combination of these two kinds of primitives on top of LLVM IR form HPVM-HDC IR. Lowering HDC++ into HPVM-HDC IR is simple. HPVM already has infrastructure for transforming opaque function calls in Hetero-C++ into LLVM intrinsics; we use this functionality to lower HDC primitives, which are opaque function calls in C++, to LLVM intrinsics. These HDC primitives become intrinsics inline the HPVM IR node functions.
HDC primitive intrinsic functions in HPVM-HDC IR can be lowered in two different ways, depending on the targeted device. They may be lowered by translating intrinsics into HPVM IR subgraphs representing the underlying computation of those HDC operation. As an example, consider compiling a Hamming distance primitive (Listing 3). HPVM-HDC can lower the primitive into a two deep loop nest, where the outer loop iterates over the classes dimension and the inner loop iterates over the hypervector dimension (Listing 4). HPVM represents parallel loop iterations as a set of dynamic instances of a node function, each with a unique ID corresponding to the loop index. In this example, the outer loop is known to be parallel, and is outlined into a separate function. The parallel loop iteration is identified by first obtaining the handle to the node in the DFG (line 6) and then querying the dynamic instance ID (line 7). Additional annotations are emitted on line 5 specifying data and classes as input edges and result as an input and output edge. In this scenario, HDC primitives are lowered into generic HPVM IR before the compiler backends are invoked.
@@
HDC primitives may instead be lowered by translating the intrinsics into calls to CUDA kernels, cuBLAS (cub, [n. d.]) calls, Thrust (thr, [n. d.]) calls, or calls to an accelerators’ functional interface. These primitives are not lowered into generic HPVM IR. Instead, the compiler backends will generate the appropriate code from the HDC primitives directly. HPVM-HDC decides which lowering strategy to take based on which backend is invoked by the user (see Section 4.3).
@@ 4.2. HDC Approximation Optimizations
To reduce the burden on the programmer in performing such rewrites, we implement an automatic binarization inter-procedural analysis and transformation described in Algorithm  1. The algorithm performs a taint-analysis to identify which HD operations and allocations are tainted by hdc_sign and then subsequently rewrites them to operate on the reduced element-bitwidth. The reduced bitwidth representation is a parameter of this transformation. For a single bit-representation, the lowering of HDC primitives are handled using bitvector logical operations. Any HDC primitives requiring additional allocation updates are binarized accordingly. By default, automatic binarization only binarizes the results of reducing operations such as matmul, cossim_similarity, and hamming_distance, and binarizes both the inputs and outputs of element-wise HDC operations. More aggressive binarization can be toggled by specifying another command-line option which reduces precision for the inputs of reducing HDC operations.
@@
Reduction Perforation. Reduction operators are common primitives in HDC algorithms—matrix multiplication is used in random projection encoding, and hamming distance and cosine similarity are used in HDC inference algorithms that identify which representative hypervector a query hypervector is most similar to. Due to the error resilience of HDC, it may be sufficient to perform these operations approximately by skipping elements along the reduction axis. We provide application developers a mechanism for controlling such approximations by exposing an HD reduction perforation directive, red_perf. The reduction computation may be approximated by performing either a segmented reduction or a strided reduction (Sidiroglou-Douskos et al., 2011). In the former, a contiguous sub-range of elements in hypervectors are used to calculate the reduction, whereas in the latter a strided access of elements in hypervectors are used. The parameters for these approximations are exposed to the application developer in their HDC++ source code.
@@
These two optimizations are fundamentally approximate. Automatic binarization reduces element bitwidth which reduces data-movement between the host and device and also improves throughput by using hardware efficient bit-wise operations. Reduction perforation approximates reduction operators between hypervectors which improves performance by reducing the total number of computations which need to be performed. These are optional transformations in HPVM-HDC; we study their implementation and empirical effects, but do not tackle automatic selection of approximations. However, they are easily explorable by an application developer; automatic binarization is exposed as compiler flags and reduction perforation is exposed as an optional HDC++ primitive that doesn’t affect algorithm code. They are applicable on the CPU and GPU, since these devices have the flexibility to execute HDC operations with different element types and loop iteration spaces. However, they are not applicable on the HDC accelerators, since these devices do not support these approximations.
@@ 4.3. Back End Code Generation
Listing 6: HPVM-HDC generated code for performing HDC training and inference on the HD Digital ASIC.
@@ 5.1. Methodology
Hardware targets. The evaluation setup for CPU and GPU consists of an Intel(R) Xeon(R) Silver 4216 CPU with 16 cores, 32 threads, and 192GB of memory and an NVIDIA GeForce RTX 2080 Ti GPU with 11GB of memory. We execute the HD-Digital ASIC as an accelerator attached to an ARM core as the host through a custom FPGA-based interface. We execute the ReRAM accelerator simulator on the Intel machine—the simulator calculates an estimate of the latency of a taped out ReRAM accelerator. For comparison with the HDC accelerators, we also execute HDC++ applications on an NVIDIA Jetson AGX Orin 64GB board with an Ampere GPU with 2048 CUDA cores and 64 tensor cores—this comparison is meant to be more representative of hardware available at the edge, which is the target use case of these HDC accelerators. Note that the Jetson AGX Orin is a moderately powerful GPU relative to other edge-class GPUs.
@@
Applications: The applications are implemented in HDC++ and summarized in Table 2, along with the computational stages in these applications. HD-Classification and HD-Clustering represent common units of computation in HDC applications, while HyperOMS, RelHD, and HD-Hashtable are three existing HDC applications from prior work. Our evaluated workloads capture most of the core computations present in other HDC applications, such as HyperSpec, SpecHD, HDBind, HyperGen, and RapidOMS (Xu et al., 2023b; Pinge et al., 2023; Jones et al., 2023; Xu et al., 2024; Pinge et al., 2024). HD-Classification and HD-Clustering codes were developed and tuned as part of (Zhang et al., 2023; Imani et al., 2019a). We used the reference implementations available for Rel-HD (Kang et al., 2022d) and HyperOMS (Kang et al., 2022c). The HD-Hashtable code is a modified version of (Zou et al., 2022), optimized for using hash tables. All five applications were developed by domain experts in prior work, and were originally hand-tuned for at least one hardware target. We evaluate HD-Classification and HD-Clustering on the Isolet (Cole and Fanty, 1994) dataset of spoken letters, HyperOMS on the Yeast and Human Spectral Library dataset and the iPRG2012 dataset (Wang et al., 2018; Chalkley et al., 2014), RelHD on the Cora dataset (McCallum, 2000) of categorized scientific papers, and HD-Hashtable on Genomics dataset adapted from (Bankevich et al., 2022).
@@ 5.2. Performance
To evaluate HPVM-HDC’s performance across architectures, we evaluate the aforementioned applications on the four hardware targets listed above. We evaluate HD-Classification and HD-Clustering on all 4 hardware targets, and HyperOMS and RelHD on the CPU and GPU only. The HDC accelerators expose coarse-grain computations for random projection encoding, training, and inference. HD-Classification maps efficiently to all three of these operations, while HD-Clustering maps to random projection encoding and inference. The other three applications do not map to these particular coarse-grained operations; we are therefore unable to execute those three applications on the HDC custom accelerators.
Figure 5 shows the relative speedup in wall clock time for each application on CPU and GPU when compiled with HPVM-HDC, compared to the baseline device-specific codes for these devices. HyperOMS lacks a CPU version we can use as the baseline—we omit the relative speedup bar in that case. In contrast, all of the applications can be compiled for and run on CPUs and GPUs using HPVM-HDC. HPVM-HDC achieves significant speed-ups over several baseline codes across architectures without using approximation-based optimizations. This advantage is of course a function of the specific implementation and tuning choices of those baselines and not a fundamental advantage of HPVM-HDC. The key contribution of HPVM-HDC is achieving competitive performance across applications and architectures while presenting a high-level, target-agnostic programming interface in HDC++.
@@
Performance on GPU. For GPUs, HPVM-HDC lowers HDC primitives into calls to CUDA kernels, cuBLAS functions, or Thrust functions, and lowers generic parallel HPVM IR to OpenCL using HPVM’s GPU backend. The only application that is slower when written in HDC++ is HyperOMS—the bottleneck portion of HyperOMS in our experiments is level ID encoding. In our HDC++ implementation, this is written using generic Hetero-C++ parallel constructs which generate OpenCL using HPVM’s GPU backend. In contrast, the baseline implementation uses an optimized CUDA kernel, including the use of warp-level primitives that HPVM cannot generate. Similar to the CPU compiled benchmarks, HD-Hashtable’s GPU compilation is interpreted in Python using CuPy, whereas HPVM-HDC generates a compiled executable. HD-Classification, HD-Clustering, HyperOMS and RelHD provide a CUDA C++ baseline implementation which present meaningful performance insights. Across the GPU compiled HPVM-HDC applications for which the reference implementation is in comparable CUDA-C++, we observe the largest performance improvements in HD-Classification with an relative improvement of 1.5x. For these 4 benchmarks, HPVM-HDC achieves a geomean 1.17x relative speedup. When including the HD-HashTable GPU performance for HPVM-HDC, we achieve a geomean speedup of 1.5x.
@@
Performance on HDC Accelerators. Figure 6 compares the performance of HD-Classification and HD-Clustering when compiling HDC++ applications for the HDC accelerators and for an NVIDIA Jetson AGX Orin board. Due to fabrication cost constraints, the digital ASIC and its ARM host CPU only communicate at approximately 10 kbps. Additionally, the ReRAM simulator only models the performance of the ReRAM accelerator, and not of a system containing the accelerator and a host. Thus, we measure the ”device-only” performance for the accelerators and the Jetson Orin, which corresponds to just the HDC primitive code in the applications.
We compare running HD-Classification and HD-Clustering on HDC accelerators against running on an NVIDIA Jetson AGX Orin board, since there is no baseline implementation targeting the accelerators for comparison—no prior work has run a whole HDC application on either of these accelerators. Additionally, while these accelerators were designed for classification  (Yang et al., 2024; Xu et al., 2023a), we can accelerate the computationally intensive portions of HD-Clustering by using stage-based HDC++ primitives. The NVIDIA Jetson AGX Orin board is representative of GPU-based compute available in edge environments—the HDC accelerators are also designed for use at the edge. We see that the accelerators deliver speed-ups for both HD-Classification and HD-Clustering compared against execution on a GPU-based system. The larger speedup in the case of HD-Classification is because this application spends most of its time training a set of representative hypervectors per class, which the HDC devices are specifically designed to accelerate. The ReRAM accelerator is estimated to be faster than the digital ASIC because the underlying processing-in-memory technology, resistive RAM, provides high density on-chip storage with fast read speeds.

### v3

@@ 3.1. High-Level Stage Primitives
In addition to describing individual HDC operations, HDC++ enables describing three commonly used HDC algorithmic stages: encoding_loop, training_loop, and inference_loop. Listing 2 illustrates expressing an HDC inference loop using the corresponding primitive. There are two advantages of providing such high-level primitives. First, it enables HPVM-HDC to target HDC accelerators, which provide coarse-grain algorithm acceleration, in contrast to acceleration of finer-grain HDC operations. Second, the use of loop intrinsics for the key algorithmic stages implemented by the HDC accelerators allows HPVM-HDC to elide data movements to and from the accelerators. Crucially, these high-level stage primitives are composable with more granular primitives in HDC++—this allows us to express certain applications that map partially, but not fully to HDC accelerators, such as HD-Clustering. By writing HD-Clustering in HDC++, we are able to map the computationally intensive part of clustering, HDC inference, to HDC accelerators, while performing more ancillary tasks, such as cluster updates and the initial generation of the random projection matrix, on the CPU or GPU.
The algorithmic stage primitives in HDC++ each take as input an ”implementation” function. This function implements encoding, training, or inference using more granular HDC primitives, and is executed when targeting CPUs or GPUs, rather than HDC accelerators. This is because while HDC accelerators implement specific encoding, training, and inference algorithms, CPUs and GPUs can be programmed to implement a variety of algorithms—it is up to the application developer to choose concrete versions of these algorithms for these targets.
1__hetero_hdc_inference_loop(InferenceFunction, QueriesHM, ClassHM, InferenceResults);
Listing 2: HDC++ example of an inference loop. QueriesHM are input vectors to classify, ClassHM contains representative hypervectors per class, InferenceResults is the destination pointer to store the results in, and InferenceFunction is a function implementing the inference algorithm using HDC primitives.
@@ 4.1. The HPVM-HDC IR
HDC++ application code is compiled into an IR that contains HDC primitives as well as generic parallel task and loop primitives. The HDC primitives are represented as LLVM intrinsic functions which represent the same operations as the HDC++ primitives listed in Table 1. We use HPVM IR primitives to represent task and data parallelism. The combination of these two kinds of primitives on top of LLVM IR form HPVM-HDC IR. Lowering HDC++ into HPVM-HDC IR is simple. HPVM already has infrastructure for transforming opaque function calls in Hetero-C++ into LLVM intrinsics; we use this functionality to lower HDC primitives, which are opaque function calls in C++, to LLVM intrinsics. These HDC primitives become intrinsics inside the HPVM IR node functions.
HDC primitive intrinsic functions in HPVM-HDC IR can be lowered in two different ways, depending on the targeted device. They may be lowered by translating intrinsics into HPVM IR subgraphs representing the underlying computation of those HDC operations. As an example, consider compiling a Hamming distance primitive (Listing 3). HPVM-HDC can lower the primitive into a two deep loop nest, where the outer loop iterates over the classes dimension and the inner loop iterates over the hypervector dimension (Listing 4). HPVM represents parallel loop iterations as a set of dynamic instances of a node function, each with a unique ID corresponding to the loop index. In this example, the outer loop is known to be parallel, and is outlined into a separate function. The parallel loop iteration is identified by first obtaining the handle to the node in the DFG (line 6) and then querying the dynamic instance ID (line 7). Additional annotations are emitted on line 5 specifying data and classes as input edges and result as an input and output edge. In this scenario, HDC primitives are lowered into generic HPVM IR before the compiler backends are invoked.
@@
Alternatively, HDC primitives may instead be lowered by translating the intrinsics into calls to CUDA kernels, cuBLAS (cub, [n. d.]) calls, Thrust (thr, [n. d.]) calls, or calls to an accelerators’ functional interface. These primitives are not lowered into generic HPVM IR. Instead, the compiler backends will generate the appropriate code from the HDC primitives directly. HPVM-HDC decides which lowering strategy to take based on which backend is invoked by the user (see Section 4.3).
@@ 4.2. HDC Approximation Optimizations
To reduce the burden on the programmer in performing such rewrites, we implement an automatic binarization inter-procedural analysis and transformation described in Algorithm  1. The algorithm performs a taint-analysis to identify which HDC operations and allocations are tainted by hdc_sign and then subsequently rewrites them to operate on the reduced element-bitwidth. The reduced bitwidth representation is a parameter of this transformation. For a single bit-representation, the lowering of HDC primitives are handled using bitvector logical operations. Any HDC primitives requiring additional allocation updates are binarized accordingly. By default, automatic binarization only binarizes the results of reducing operations such as matmul, cossim_similarity, and hamming_distance, and binarizes both the inputs and outputs of element-wise HDC operations. More aggressive binarization can be toggled by specifying another command-line option which reduces precision for the inputs of reducing HDC operations.
@@
Reduction Perforation. Reduction operators are common primitives in HDC algorithms—matrix multiplication is used in random projection encoding, and hamming distance and cosine similarity are used in inference algorithms that identify which representative hypervector a query hypervector is most similar to. Due to the error resilience of HDC, it may be sufficient to perform these operations approximately by skipping elements along the reduction axis. We provide application developers a mechanism for controlling such approximations by exposing an HDC reduction perforation directive, red_perf. The reduction computation may be approximated by performing either a segmented reduction or a strided reduction (Sidiroglou-Douskos et al., 2011). In the former, a contiguous sub-range of elements in hypervectors are used to calculate the reduction, whereas in the latter a strided access of elements in hypervectors are used. The parameters for these approximations are exposed to the application developer in their HDC++ source code.
@@
1hypervector<2048> A, B = ...;
2// Calculate hamming distance in the sub-range [0, 1024)
3auto dst_segmented = __hetero_hdc_hamming_distance(A, B);
4__hetero_hdc_red_perf(dst_segmented, 0, 1024, 1);
5// Calculate hamming distance for every other element
6auto dst_strided = __hetero_hdc_hamming_distance(A, B);
7__hetero_hdc_red_perf(dst_strided, 0, 2048, 2);
8// Hamming distance for every other element in [0, 1024)
9auto dst_both = __hetero_hdc_hamming_distance(A, B);
10__hetero_hdc_red_perf(dst_both, 0, 1024, 2);
@@
These two optimizations are fundamentally approximate. Automatic binarization reduces element bitwidth which reduces data-movement between the host and device and also improves throughput by using hardware efficient bit-wise operations. Reduction perforation approximates reduction operators between hypervectors which improves performance by reducing the total number of computations which need to be performed. These are optional transformations in HPVM-HDC; we study their implementation and empirical effects, but do not tackle automatic selection of approximations. However, they are easily explorable by an application developer; automatic binarization is exposed as a compiler flag and reduction perforation is exposed as an optional HDC++ primitive that doesn’t affect algorithm code. They are applicable on the CPU and GPU, since these devices have the flexibility to execute HDC operations with different element types and loop iteration spaces. However, they are not applicable on the HDC accelerators, since these devices do not support these approximations.
@@ 4.3. Back End Code Generation
1config_t config = ...;
2initialize_device(&config);
3allocate_base_mem(random_projection);
4allocate_class_mem(classes);
5for (int n = 0; n < EPOCHS; ++n)
6 for (int i = 0; i < N_TRAIN; ++i)
7 allocate_feature_mem(train_inputs[i]);
8 execute_retrain(train_labels[i]);
9read_class_mem(classes);
10allocate_base_mem(random_projection);
11allocate_class_mem(classes);
12for (int i = 0; i < N_TEST; ++i)
13 allocate_feature_mem(infer_inputs[i]);
14 infer_labels[i] = execute_inference();
Listing 6: HPVM-HDC generated code for performing HDC training and inference on the digital HDC ASIC.
@@ 5.1. Methodology
Hardware targets. The evaluation setup for CPU and GPU consists of an Intel(R) Xeon(R) Silver 4216 CPU with 16 cores, 32 threads, and 192GB of memory and an NVIDIA GeForce RTX 2080 Ti GPU with 11GB of memory. We execute the digital HDC ASIC as an accelerator attached to an ARM core as the host through a custom FPGA-based interface. We execute the ReRAM HDC accelerator simulator on the Intel machine—the simulator calculates an estimate of the latency of a taped out ReRAM accelerator. For comparison with the HDC accelerators, we also execute HDC++ applications on an NVIDIA Jetson AGX Orin 64GB board with an Ampere GPU with 2048 CUDA cores and 64 tensor cores—this comparison is meant to be more representative of hardware available at the edge, which is the target use case of these HDC accelerators (the Jetson AGX Orin is a moderately powerful GPU relative to other edge-class GPUs).
@@
Applications: We implement five applications, summarized in Table 2, in HDC++. All five applications were developed by domain experts in prior work. HD-Classification and HD-Clustering represent common computational kernels in HDC applications, while HyperOMS, RelHD, and HD-Hashtable are three existing HDC applications from prior work. Our evaluated workloads capture most of the core computations present in other HDC applications, such as HyperSpec (Xu et al., 2023b), SpecHD (Pinge et al., 2023), HDBind (Jones et al., 2023), HyperGen (Xu et al., 2024), and RapidOMS (Pinge et al., 2024). HD-Classification and HD-Clustering codes were developed and tuned as part of (Zhang et al., 2023) and (Imani et al., 2019a), respectively. We used the reference implementations available for Rel-HD (Kang et al., 2022d) (CPU and GPU) and HyperOMS (Kang et al., 2022c) (just GPU). The HD-Hashtable code is a modified version of (Zou et al., 2022), optimized for using hash tables. We evaluate HD-Classification and HD-Clustering on the Isolet (Cole and Fanty, 1994) dataset of spoken letters, HyperOMS on the Yeast and Human Spectral Library dataset and the iPRG2012 dataset (Wang et al., 2018; Chalkley et al., 2014), RelHD on the Cora dataset (McCallum, 2000) of categorized scientific papers, and HD-Hashtable on Genomics dataset adapted from (Bankevich et al., 2022). When targeting the GPU, these baseline codes are often written in CUDA C++. While unportable, this enables the use of specific device features that improve performance, such as tensor cores and warp voting functions.
@@ 5.2. Performance
To evaluate HPVM-HDC’s performance across architectures, we compile and execute the aforementioned applications on the four hardware targets listed above. We evaluate HD-Classification and HD-Clustering on all 4 hardware targets, and HyperOMS and RelHD on the CPU and GPU only. The HDC accelerators expose coarse-grain computations for random projection encoding, training, and inference. HD-Classification maps efficiently to all three of these operations, while HD-Clustering maps to random projection encoding and inference. The other three applications do not map to these particular coarse-grained operations; we are therefore unable to execute those three applications on the HDC custom accelerators.
Figure 5 shows the relative speedup in wall clock time for each application on CPU and GPU when compiled with HPVM-HDC, compared to the baseline device-specific codes for these devices from prior work. HyperOMS lacks a CPU version we can use as the baseline—we omit the relative speedup bar in that case. In contrast, all of the applications can be compiled for and run on CPUs and GPUs using HPVM-HDC. HPVM-HDC achieves significant speed-ups over several baseline codes across architectures without using approximation-based optimizations. However, this advantage is a function of the specific implementation and tuning choices of those baselines and not a fundamental advantage of HPVM-HDC. The key contribution of HPVM-HDC is achieving competitive performance across applications and architectures while presenting a high-level, target-agnostic programming interface in HDC++.
@@
Performance on GPU. For GPUs, HPVM-HDC lowers HDC primitives into calls to CUDA kernels, cuBLAS functions, or Thrust functions, and lowers generic parallel HPVM IR to OpenCL using HPVM’s GPU backend. The only application that is slower when written in HDC++ is HyperOMS, which is 5% slower than the baseline. The bottleneck portion of HyperOMS in our experiments is level ID encoding. In our HDC++ implementation, this is written using generic Hetero-C++ parallel constructs which generate OpenCL using HPVM’s GPU backend. In contrast, the baseline implementation uses an optimized CUDA kernel, including the use of warp-level primitives that HPVM cannot generate.
Similar to the CPU compiled benchmarks, HD-Hashtable’s GPU compilation is interpreted in Python using CuPy, whereas HPVM-HDC generates a compiled executable. HD-Classification, HD-Clustering, HyperOMS and RelHD provide a CUDA C++ baseline implementation, so the comparisons for these can be used to derive meaningful performance insights. Across the GPU compiled HPVM-HDC applications for which the reference implementation is in CUDA C++, we observe a range of 0.95x – 1.5x. The largest performance improvement (1.5x) is in HD-Classification. For these 4 benchmarks, HPVM-HDC achieves a geomean 1.17x relative speedup. This speedup is primarily the result of different tuning choices in HPVM-HDC generated code compared to the baseline codes.
@@
Performance on HDC Accelerators. Figure 6 compares the performance of HD-Classification and HD-Clustering when compiling HDC++ applications for the HDC accelerators and for an NVIDIA Jetson AGX Orin board. Due to fabrication cost constraints, the digital ASIC and its ARM host CPU only communicate at approximately 10 kbps. Additionally, the ReRAM accelerator simulator only models the performance of the ReRAM accelerator, and not of a system containing the accelerator and a host. Thus, we measure the ”device-only” performance for the accelerators and the Jetson Orin, which corresponds to just the HDC primitive code in the applications.
We compare running HD-Classification and HD-Clustering on HDC accelerators against running on an NVIDIA Jetson AGX Orin board, since there is no baseline implementation targeting the accelerators for comparison—no prior work has run a whole HDC application on either of these accelerators. Additionally, while these accelerators were designed for classification  (Yang et al., 2024; Xu et al., 2023a), we can accelerate the computationally intensive portions of HD-Clustering by using HDC++ primitives. The NVIDIA Jetson AGX Orin board is representative of GPU-based compute available in edge environments—the HDC accelerators are also designed for use at the edge. We see that the accelerators deliver speed-ups for both HD-Classification and HD-Clustering compared against execution on a GPU-based system. The larger speedup in the case of HD-Classification is because this application spends most of its time training a set of representative hypervectors per class, which the HDC devices are specifically designed to accelerate. The ReRAM accelerator is estimated to be faster than the digital ASIC because the underlying processing-in-memory technology, resistive RAM, provides high density on-chip storage with fast read speeds.